OpenStudy (anonymous):

Mona Galland recieved a year end bonus of $17,000 from her company and invested the money in an account paying 6.5%. How much additional money should she deposit in an account paying 5% so that the return in the two investments will be 6%?

jimthompson5910 (jim_thompson5910):

If you invest $17,000 at an interest rate of 6.5%, then you will earn 17000*0.065 = 1,105 dollars in interest alone This is just from the initial investment of $17,000

jimthompson5910 (jim_thompson5910):

Let x = the amount of money you additionally invest in the 5% account so if you invest x dollars at 5% interest, then you will earn 0.05x dollars in interest alone

jimthompson5910 (jim_thompson5910):

In total, you have invested 17000 + x dollars Of this total investment, you have earned a total of 1105+0.05x dollars in interest alone

jimthompson5910 (jim_thompson5910):

The ratio of interest to total invested must be 6% because this is what the problem originally states, so... (Amount earned in interest)/(Total amount invested) = Overall Yield (1105 + 0.05x)/(17000 + x) = 0.06 Now solve for x to get your answer